# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ 4.1.4. Challenges and shortcomings encountered in using GenAI
Respondents also identified several challenges arising from the use of GenAI, the most significant of which was the generation of unhelpful, unwanted, or broken code (128). Additional challenges included hallucinations (64) (such as recommending non-existent libraries or method calls), problems arising from outdated training data (53), issues arising from tools not having access to the project’s wider context (52), which could be caused by insufficient context windows (40), and an inability to use GenAI to solve complex problems (43). $R_{291}$ noted, “[A]s issues get more complex, code generation has many gaps.” $R_{148}$ elaborates: “AI can do simple code pretty well, it’s not that capable [of generating] more complex code, it’s incapable of generating well-architected code outside of extremely simple patterns.” $R_{33}$ provided an example: “In my experience, a simple Python/Django project that required some Form customizations was just not possible for any AI tool at my disposal. The logical part, the thinking, and the experience gained from years of coding real-world applications and systems design are [not comparable to] an AI code generation tool like GPT or Gemini.” The full list of challenges reported by developers is included in our replication package (ano, 2024).
Twenty-five respondents also noted that while the ability to write good prompts is essential to effectively using GenAI tools, it can present a significant challenge. $R_{303}$ reported that “AI cannot complete the majority of tasks by itself, even simple ones[,] without time spent [by the developer on] prompt engineering.” As $R_{378}$ put it, “Models need ‘hand-holding’ instructions [in order] to generate code and debug it.” But when crafting these prompts, “[q]uestions must be asked precisely,” as $R_{135}$ commented. $R_{432}$ elaborated, “Vague prompts lead to vague answers, so you have to be specific with what you want.” $R_{72}$ summarized all these sentiments: “The outputs [of GenAI tools] are only as good as the inputs.”
Writing detailed and specific prompts isn’t always an easy task, however. $R_{18}$ said, “It is hard to communicate what I want or what I’m trying to describe to the model.” $R_{83}$ echoed this concern, stating, “Sometimes I am unable to describe what it is that [I] want[,] only to find out [that] there existed vocabulary that I should have used[, but] [t]his issue seems less frequent recently compared to when [I] first started using [GenAI] for coding.” If done correctly, however, prompting was also seen as a way to overcome other limitations of the technology. $R_{240}$ said, “There is a limit to the complexity that can be generated stand[ing] alone, but a human crafting intelligent prompts can still architect complex solutions using AI.”
Finding 5: Aside from obvious problems such as possible hallucinations, GenAI for coding still suffers from challenges related to lack of suitable contextual information available to the model at inference time, the inability for such models to solve complex, non-trivial tasks such as highly-specific functionality which conforms to a given architecture, and the need to provide good, well-designed prompts.
@@ 4.2.1. AI-generated code review and compliance
Given the potential challenges associated with the use of GenAI, mitigation strategies relating to vulnerability and compliance analysis become increasingly important. In our follow-up interviews, we asked participants additional questions regarding their workflows. (Section 3.5 provides demographic information on participants, including development experience background.)
As $RI_{118}$ noted, when code is not generated by the developer, review can be a challenge: “[T]hat’s another big part of it is just going through line by line, but it’s like a more passive mindset, which I find a bit challenging compared to actually writing the code, so it can be harder to actually spot what’s wrong with it. So the tests become really important.” This is particularly true with respect to edge cases, “because the tools are actually quite good, but unimaginative and have no domain, [they] have no context.”
Some developers reported an overall review but not focusing on vulnerabilities. $RI_{190}$ said, “[I don’t do] any checks for licensing or vulnerabilities, but usually I read the code before I commit it and it’s also reviewed in a pull request.” $RI_{39}$ offered a similar approach: “I think it’s mostly that I read the code and try and evaluate first of all how dangerous it is to run.”
@@
Others believed that AI itself could assist with reviewing AI-generated code. $RI_{149}$ told us that they “ask the GenAI itself to review the code, perhaps using a different model […] [and] ask it to give me gotchas or things to look out for.” $RI_{31}$ tentatively agreed but acknowledged the challenges: “I’ve thought about this […,] and [if] it’s something where the answer is using machine learning to check whether or not the code that you’ve used appears, and would put you at legal risk, […] you’re relying on a large language model to check whether your large language model puts you at legal risk.”
@@
Some developers, however, employed manual checks, particularly for license compliance. $RI_{516}$ described a process that involved “search[ing] on GitHub because that’s the largest repository, but […] also check[ing] inputs in search engines, finding other places like GitLab or CodeBerg or Bitbucket or other repositories. And if the code is in a visible repository, it should show up.” $RI_{431}$ would “try to do my own research on websites such as Stack Overflow or GitHub.” $RI_{431}$ emphasized the manual nature of their compliance review: “I don’t really have any tool [to verify provenance], so that’s why I’m doing this step of going to a website such as Stack Overflow and trying to see if the code which has been given to me by the model has already been proposed somewhere else on the Internet.” But $RI_{516}$ noted that these manual processes are limited, because “if it’s not open source, I probably have no way to or means to check it because closed source is not generally available […] in a search or anything like that.”
@@
Some developers stated that they avoided compliance checks on AI-generated code for fear of opening themselves or their organizations to liability. $RI_{39}$ explained, “I have gotten advice from IP lawyers at a previous job […] to not look at patents for anything that I’m working on. Because if there is evidence that I have seen a patent that we then infringe on, then it’s willful infringement, which can mean triple damages or something like that. […] I could imagine a case where it might make sense to not track down which lines of code were generated by a LLM, because then you might have less of a willful infringement case or something like that.”
@@ 4.2.2. Documentation of GenAI usage
Figure 7. Developer awareness of documentation processes for GenAI-related tasks within their organizations
@@
One might expect that organizations that permit the use of GenAI for code development would ensure that developers were aware of processes to document its use. However, of the 445 developers indicating that their organization (partially) permitted the use of GenAI, only 83 (18.7%) indicated that they were aware of such a process, and of the remaining group, 312 (70.1%) said that there was no process and another 50 (11.2%) were unsure if a process existed. $R_{325}$ offered one view of why a process was unnecessary: “I did not use anything that cannot be googled anyway—but we never thought about having to document or question code we found through Google.” Whether this means that many organizations have no process for documenting the use of GenAI or that at least some developers incorrectly believe that no documentation process exists at their organization, the result may be that developers do not have a full understanding of the possible legal risks they or their organizations face.
@@
Finding 7: Most developers in our study (70.1%) were unaware of any process for documenting GenAI usage within their organization. To the extent that organizations have processes in place for documenting the use of GenAI, they may not be sufficiently educating their employees about those processes.
@@
We further asked the 83 respondents who reported that their companies had processes to document their AI usage about what information was collected during the process. The majority of responses (49, 59.0%) indicated that documentation consisted of noting the use of an AI tool, with 36 respondents (43.4%) indicating that the prompt was also documented. In 27 cases (32.5%), the entire interaction with the code generation tool was documented. Other documentation strategies included keeping system/configuration records (3), summarizing the purpose of using AI tooling (2), providing links to chats (1), documenting known risks associated with the use of AI (1), storing the AI agent’s memory (1), and relying on version control (1).
@@
Finding 8: The most common documentation process among respondents was to note the fact of GenAI usage and/or to document the prompt used to acquire the generation. However, nearly a third of respondents (32.5%) indicated documenting the whole interaction, potentially leading to reproducibility.
@@
Some respondents indicated that one motivating factor behind documentation, at least for developers, was to explain the cause of a potential failure rather than to provide provenance. $R_{190}$ noted, “I usually write a comment ‘written by [C]hatGPT’ or ‘partly written by [C]hatGPT’ to denote that if it breaks, I would maybe have written it better but used [C]hatGPT to speed up the process.” $RI_{431}$ echoed this sentiment during a follow-up interview: “But when I do use artificial intelligence to generate code […], I will try to notify in a code comment or in a comment inside the documentation that this part needs to be rewritten or at least tested […] [so that we can] change anything that wouldn’t work in our current case.” $RI_{118}$ suggested that the typical training at their place of employment does not anticipate this type of activity: “[Developers are] allowed to start using the tools and there’s guidance on where and when they can’t, but not in terms of attribution or [adding] warning[s that say,] ‘Hey, here be code generated monsters.”’
@@ 4.2.3. Review of the Terms of Service for AI Tools
The Terms of Service (ToS) for GenAI tools lay out how developers may interact with models by describing permissible and impermissible behaviors. These ToS may contain language that prevents the use of the service for any illegal activity, forbids the use of model output as training data for another, competing model, or, for open-weight models, provide stipulations on what is allowed when fine-tuning. The ToS may also include stipulations that data provided as prompts by a user may be used as training data for future models. In short, it is important for both users and model fine-tuners to understand the terms that they are agreeing to before using a model.
@@
Developers’ views on the ToS of GenAI tools varied widely. A little over a third of respondents who said that they incorporated AI into their workflow (177, 35.6%) indicated that they never read terms in full but may skim them looking for relevant or problematic provisions. A similar percentage of respondents (141, 28.4%) stated that they sometimes read the ToS in full, depending on the tool they are using. An additional 127 respondents (25.6%) claimed that they never read the ToS for the tools they use. The smallest group, 52 participants (10.5%), told us that they always read the ToS in full.
In follow-up interviews, respondents who tended not to read the ToS provided their rationale. For one participant, the decision was based on other priorities. $RI_{190}$ explained, “[As to] why I don’t read it, it’s because I don’t have time. We have features to ship, and I work at a startup, so we need to find this Holy Grail, which is product market fit. […] It doesn’t feel like reading the Terms of Service will help us, so I don’t do it.” $RI_{431}$ indicated that the main reason they do not read the ToS is “quite often the length of the text.”
@@
Still other respondents indicated that they relied on others to read the documents and alert them to relevant provisions. Respondents who worked within organizations stated that they trusted that their organization’s legal department would take responsibilty for compliance. $RI_{39}$ stated, “I personally don’t worry about [licensing issues related to generated code] because the legal department where I work has already authorized [the] use of [the LLMs], and it’s kind of their problem from my perspective.”
@@
Others noted that they relied on discussions on blogs and social media to identify problematic provisions. $RI_{190}$ stated, “If it would blow up on the internet that the terms are bad, then I would kind of hear about it.” $RI_{31}$ indicated they have typically “already heard about pros and cons” of using certain models and described one rule they were “trying to stick by”: “I know that some models and services just copy and paste what’s already been created […] It’s stealing work from someone. I don’t really agree with that philosophy.” $RI_{31}$ also used models’ ToS and social media to identify models that purportedly trained on unauthorized data: “You can review their Terms of Service, and look at their blogs and so on. The other piece is just honestly checking Reddit [and] seeing what other people are talking about.” $RI_{516}$ referred to the website Terms of Service; Didn’t Read (https://tosdr.org) which “[does] a summary of the main points of each Terms of Service and privacy policy on the services that have already been analyzed. It’s analyzed by a community […]. The main points of concern […] are marked in red, so you need to watch out for them.”
@@ 4.2.4. Copyright/legal training
Of the 574 respondents, only 68 (11.9%) indicated that they had any formal training in copyright law and/or the legal implications of using code generation models. Of these, 28 (41.2%) sought the training out on their own. Sixteen respondents (23.5%) indicated that the training was provided by their employer, and 24 (35.3%) described their training as some combination of self-initiated and employer-provided. (Some respondents interpreted the question to refer to GenAI training more generally rather than training that focused on legal implications.)
@@
Finding 10: Relatively few respondents (11.9%) reported having undergone any formal training on the legal implications of GenAI. Some struggled to find suitable available resources.
Table 5. Main findings for RQ1: Use of GenAI for software development
4.1 Uses, benefits, and challenges of GenAI
4.1.1 GenAI usage in Software Development
F1
While the vast majority of developers surveyed (89.7%) were using GenAI tools, there were many
reasons why developers might not incorporate GenAI into their development workflows, including
perceived legal issues, personal preference, information security concerns, and limitations placed
by their employers.
F2
Of those respondents who reported working at organizations that disallowed the use of GenAI, many
indicated that they believed the restriction was based on fear of proprietary information leakage and
the potential legal challenges such usage might bring. Some developers also perceived that their
employers were slow to adopt new technology.
4.1.2 GenAI tools used by developers
F3
The ecosystem of GenAI tools used by developers was varied and extensive, comprising both
open-source/open-weight and proprietary solutions. Developers primarily made use of web-based
tools and those that could be integrated into their IDEs.
4.1.3 Benefits derived from GenAI usage
F4
The benefits developers reported from using GenAI tools for coding go beyond increasing productivity
and avoiding repetitive, boring tasks. They also included understanding code better and improved code
quality and code optimization, which some human developers might struggle to achieve.
4.1.4 Challenges and shortcomings encountered in using GenAI
F5
Aside from obvious problems such as possible hallucinations, GenAI for coding still suffers from
challenges related to lack of suitable contextual information available to the model at inference time,
the inability for such models to solve complex, non-trivial tasks such as highly-specific functionality
which conforms to a given architecture, and the need to provide good, well-designed prompts.
4.2 Current practices for documentation and compliance
4.2.1 AI-generated code review and compliance
F6
Among respondents, GenAI code review is typically done manually or by using other GenAI to check
for code correctness and compliance. Some developers are concerned that the process of conducting a
review may expose them or their employers to liability.
4.2.2 Documentation of GenAI usage
F7
Most developers in our study (70.1%) were unaware of any process for documenting GenAI usage within
their organization. To the extent that organizations have processes in place for documenting the use of
GenAI, they may not be sufficiently educating their employees about those processes.
F8
The most common documentation process among respondents was to note the fact of GenAI usage and/or
to document the prompt used to acquire the generation. However, nearly a third of respondents (32.5%)
indicated documenting the whole interaction, potentially leading to reproducibility.
4.2.3 Review of Terms of Service
F9
Even though some provisions in a tool’s Terms of Service may be important to developers and their work,
developers do not always thoroughly read the Terms of Service for the tools they use, choosing instead to
skim the document, assume that their organization has read the document, or rely on information and
validation from others in the development community.
4.2.4 Copyright/legal training
F10
Relatively few respondents (11.9%) reported having undergone any formal training on the legal
implications of GenAI. Some struggled to find suitable available resources.
@@ 5. RQ2: Developers’ perceptions of licensing and copyright issues emerging from the use of Generative AI
RQ2 sought to determine developer’s beliefs and understandings regarding the licensing and copyright issues that arise from the use of GenAI tools. Here we explore developers’ thoughts on current litigation and the prospect of regulation, sentiments regarding the use of their own code in model training data, opinions on the ownership of GenAI output and prompts, and lastly, the perceived risk of copyright infringment. A summary of key findings can be found in Table 6.
@@ 5.1. Perceptions about litigation and regulation
We asked respondents whether they believed that litigation would result in limitation on the use of GenAI for software development. (Figures 10 and 11 show respondents’ answers to nine five-point Likert-scale (Oppenheim, 2000) questions.) Of the 554 respondents who indicated a familiarity with AI tools, 205 (37.0%) somewhat agreed with the view that current litigation will result in limitations placed on the use of GenAI models for software development, with 76 (13.7%) respondents strongly agreeing. An additional 109 (19.7%) respondents were neutral (perhaps reflective of the legal uncertainties at play), and the third largest group of 104 developers (18.8%) somewhat disagreed. The smallest group, 60 developers (10.8%), strongly disagreed.
While more respondents believed that litigation would result in limitations on the use of GenAI, developers who believed that litigation would not result in limitations were more vocal about their views in their responses to open-ended follow-up questions. Indeed, many respondents indicated that technology would outpace attempts to regulate it, whether through litigation or otherwise. $R_{98}$ commented, “I don’t see any limitations coming due to ongoing litigation—yes, it may cause issues for OpenAI and StabilityAI, but you can’t put the genie back in the lamp.” $R_{507}$ further elaborated: “Pandora’s box has been opened. Even if the current models were legally limited, good luck proving that certain code has been used as training data, and good luck stopping people from running Llama instances locally.” $R_{54}$ acknowledged the existence of legal issues but believed that “the momentum that’s already been built by OpenAI and all the hype around the tools means there will not be the political capital to get any restrictive regulations in place.” $R_{250}$ said, “Generative AI is used by almost all software developers at this point. That ship has sailed. Lawyers should give up on even trying to figure this out.”
@@
During a follow-up interview, $RI_{118}$ mused on a potential regulatory body: “No one has solved regulatory capture, so […] I don’t know how to solve that one. It would probably have to be like a NIST [National Institute of Standards and Technology] or some kind of body like that. […] Hopefully [it] isn’t run by presidential appointment, so the datasets are not decided by the politics of the day [and not by] an industry rubber stamp body because that doesn’t help either. […] So it would probably have to be some standards organization.”
Finally, some respondents conveyed the view that it would not be technology that changed to adapt to the law but rather the law that would change to adapt to technology. $R_{376}$ commented, “The result […] will probably not be limitations on the technology since open-source is rapidly developing and cannot be genuinely regulated in this fashion. Rather, intellectual property, artistic rights, etc. will need to be recalibrated.” $R_{79}$ echoed this sentiment: “The old style copyright will not make any sense as we know it as we are entering the bumpy road to abundance.” $R_{8}$ similarly put the burden on the law to adapt: “The legal side will be resolved by lawyers when they catch up with what these tools actually do.” And $R_{182}$ offered an aspirational approach, stating that they “don’t think it is wise to approach the question of copyright of AI-generated/assisted code as a matter of whether the generated code is or is not copyrightable based primarily on what has been settled. Rather, we should think about what we want it to be, keeping in mind the benefits copyright is intended to bring to humanity.”
@@
5.2. Using one’s own code in models’ training data or as output
If regulation of GenAI occurs, whether through legislation or litigation, it should ideally take account of the perceptions and practices of stakeholders, including software developers. Our study, therefore, sought to learn respondents’ views regarding the use of code as training data for generative models.

### v3

@@ 4.1.4. Challenges and shortcomings encountered in using GenAI
Respondents also identified several challenges arising from using GenAI, the most significant of which was the generation of unhelpful, unwanted, or broken code (128). Additional challenges included hallucinations (64) (such as recommending non-existent libraries or method calls), problems arising from outdated training data (53), issues arising from tools not having access to the project’s wider context (52), which could be caused by insufficient context windows (40), and an inability to use GenAI to solve complex problems (43). $R_{148}$ elaborated: “AI can do simple code pretty well, it’s not that capable [of generating] more complex code, it’s incapable of generating well-architected code outside of extremely simple patterns.” $R_{33}$ provided an example: “In my experience, a simple Python/Django project that required some form customizations was just not possible for any AI tool at my disposal. The logical part, the thinking, and the experience gained from years of coding real-world applications and systems design are [not comparable to] an AI code generation tool like GPT or Gemini.” The full list of challenges reported by developers is included in our replication package (ano, 2024).
Twenty-five respondents noted that the ability to write good prompts is essential to effectively using GenAI tools. $R_{303}$ reported that “AI cannot complete the majority of tasks by itself, even simple ones[,] without time spent [by the developer on] prompt engineering.”
Writing effective prompts is not always an easy task, however. $R_{18}$ said, “It is hard to communicate what I want or what I’m trying to describe to the model.” If done correctly, however, prompting was also seen as a way to overcome other limitations of the technology. $R_{240}$ said, “There is a limit to the complexity that can be generated stand[ing] alone, but a human crafting intelligent prompts can still architect complex solutions using AI.”
Finding 5: Aside from obvious problems such as possible hallucinations, participants reported that the use of GenAI for coding presents challenges related to the lack of suitable contextual information available to the model at inference time, the inability of such models to solve complex, non-trivial tasks, and the need to provide good, well-designed prompts.
@@ 4.2.1. AI-generated code review and compliance
Given the potential challenges associated with using GenAI, strategies to mitigate security and compliance risks become increasingly important. (Those strategies can differ depending on the environment. Section 3.5 provides demographic information on participants, including development experience background.)
@@
As $RI_{118}$ noted, when code is not generated by the developer, review can be a challenge: “[T]hat’s another big part of it is just going through line by line, but it’s like a more passive mindset, which I find a bit challenging compared to actually writing the code, so it can be harder to actually spot what’s wrong with it.” Accordingly, two interview respondents believed that AI could assist with reviewing AI-generated code. $RI_{149}$ told us that they “ask the GenAI itself to review the code, perhaps using a different model […] [and] ask it to give me gotchas or things to look out for.” $RI_{31}$ tentatively agreed but acknowledged the challenges: “I’ve thought about this […,] and [if] it’s something where the answer is using machine learning to check whether or not the code that you’ve used appears, and would put you at legal risk, […] you’re relying on a large language model to check whether your large language model puts you at legal risk.”
@@
Two interview respondents reported employing manual checks on sites like GitHub or Stack Overflow to determine the provenance of AI-generated code to assess, for example, whether they are in compliance with that code’s license. $RI_{431}$ emphasized the manual nature of their compliance review: “I don’t really have any tool [to verify provenance], so that’s why I’m doing this step of going to a website such as Stack Overflow and trying to see if the code which has been given to me by the model has already been proposed somewhere else on the Internet.” But $RI_{516}$ noted that these manual processes are limited, because “if it’s not open source, I probably have no way to or means to check it because closed source is not generally available […] in a search or anything like that.”
@@
Other developers reported an overall review that did not focus on vulnerabilities, and still others stated that they avoided compliance checks on AI-generated code altogether for fear of opening themselves or their organizations to liability. $RI_{39}$ explained, “I have gotten advice from IP lawyers at a previous job […] to not look at patents for anything that I’m working on. Because if there is evidence that I have seen a patent that we then infringe on, then it’s willful infringement, which can mean triple damages or something like that. […] I could imagine a case where it might make sense to not track down which lines of code were generated by a LLM, because then you might have less of a willful infringement case or something like that.”
@@ 4.2.2. Documentation of GenAI usage
One might expect that organizations that permit using GenAI for code development would ensure that developers were aware of processes to document its use so that the organization’s legal risk could be assessed by appropriate personnel. However, of the 445 developers indicating that their organization (partially) permitted the use of GenAI, only 83 (19%) indicated that they were aware of such a process, and of the remaining group, 312 (70%) said that there was no process and another 50 (11%) were unsure if a process existed. $R_{325}$ offered one view of why a process was unnecessary: “I did not use anything that cannot be googled anyway—but we never thought about having to document or question code we found through Google.” Whether this means that many organizations have no process for documenting the use of GenAI or that at least some developers incorrectly believe that no documentation process exists at their organization, the result may be that developers do not have a full understanding of the possible legal risks they or their organizations face (as described in Section 2.1).
@@
Finding 7: Most developers in our study (81%) reported that there was no process for documenting GenAI usage within their organization or that they were unaware of any such process. To the extent that organizations have processes in place for documenting the use of GenAI, they may not be sufficiently educating their employees about those processes.
@@
We further asked the 83 respondents who reported that their companies had processes to document their AI usage about what information was collected during the process. The majority of responses (49, 59%) indicated that documentation consisted of noting the use of an AI tool, with 36 respondents (43%) indicating that the prompt was also documented. In 27 cases (33%), the entire interaction with the code generation tool was documented. Other documentation strategies included keeping system/configuration records (3), summarizing the purpose of using AI tooling (2), providing links to chats (1), documenting known risks associated with the use of AI (1), storing the AI agent’s memory (1), and relying on version control (1).
@@
Finding 8: The most common documentation process among respondents was to note the fact of GenAI usage and/or to document the prompt used to acquire the generation. However, nearly a third of respondents (33%) indicated documenting the whole interaction, potentially leading to reproducibility.
@@
One motivating factor behind documentation was to explain the cause of a potential failure rather than to provide provenance. $R_{190}$ noted, “I usually write a comment ‘written by [C]hatGPT’ or ‘partly written by [C]hatGPT’ to denote that if it breaks, I would maybe have written it better but used [C]hatGPT to speed up the process.” $RI_{118}$ suggested, by contrast, that the typical training at their place of employment does not anticipate this type of activity: “[Developers are] allowed to start using the tools and there’s guidance on where and when they can’t, but not in terms of attribution or [adding] warning[s that say,] ‘Hey, here be code generated monsters.”’
@@ 4.2.3. Review of the Terms of Service for AI Tools
The Terms of Service (ToS) for GenAI tools lay out how developers may interact with models by describing permissible and impermissible behaviors. These ToS may contain language that prevents the use of the service for any illegal activity, forbids the use of model output as training data for another competing model, or (for open-weight models) provides stipulations on what is allowed when fine-tuning. The ToS may also include stipulations that data provided as prompts by a user may be used as training data for future models. In short, it is important for both users and model fine-tuners to understand the terms that they are agreeing to before using a model.
@@
Developers’ views on the ToS of GenAI tools varied widely. A little over a third of respondents who said they incorporated AI into their workflow (177, 36%) indicated that they never read terms in full but may skim them looking for relevant or problematic provisions. A similar percentage of respondents (141, 28%) stated that they sometimes read the ToS in full, depending on the tool they are using. An additional 127 respondents (26%) claimed that they never read the ToS for the tools they use. The smallest group, 52 participants (11%), told us that they always read the ToS in full.
In follow-up interviews, respondents who tended not to read the ToS provided their rationale. This decision was often based on factors related to time, with $RI_{431}$ citing the “length of the text,” or having other priorities, as indicated by $RI_{190}$, who described having more impactful development work to do than reading the ToS.
@@
Still other respondents indicated that they relied on others to read the documents and alert them to relevant provisions. Respondents who worked within organizations stated that they trusted that their organization’s legal department would take responsibility for compliance. $RI_{39}$ stated, “I personally don’t worry about [licensing issues related to generated code] because the legal department where I work has already authorized [the] use of [the LLMs], and it’s kind of their problem from my perspective.”
@@
Others noted that they relied on discussions on blogs and social media to identify problematic provisions. $RI_{190}$ stated, “If it would blow up on the internet that the terms are bad, then I would kind of hear about it.” $RI_{516}$ referred to the website Terms of Service; Didn’t Read (https://tosdr.org), which “[does] a summary of the main points of each Terms of Service and privacy policy on the services that have already been analyzed. It’s analyzed by a community […]. The main points of concern […] are marked in red, so you need to watch out for them.”
@@ 4.2.4. Copyright/legal training
Of the 574 respondents, only 68 (12%) indicated that they had any formal training in copyright law and/or the legal implications of using code generation models. Of these, 28 (41%) sought the training out on their own. Sixteen respondents (24%) indicated that the training was provided by their employer, and 24 (35%) described their training as some combination of self-initiated and employer-provided. (Some respondents interpreted the question to refer to GenAI training more generally rather than training that focused on legal implications.)
@@
Finding 10: Relatively few respondents (12%) reported having undergone any formal training on the legal implications of GenAI. Some struggled to find suitable available resources.
@@ 5. RQ2: Developers’ perceptions of licensing and copyright issues emerging from the use of Generative AI
RQ2 sought to determine developers’ beliefs and understandings regarding the licensing and copyright issues that arise from using GenAI tools. Here, we explore developers’ thoughts on current litigation and the prospect of regulation, sentiments regarding the use of their own code in model training data, opinions on the ownership of GenAI output and prompts, and the perceived risk of copyright infringement. This research question is answered by Findings 11-24, which are summarized in our replication package (ano, 2024).
@@ 5.1. Perceptions about litigation and regulation
We asked respondents whether they believed that litigation would result in limitations on the use of GenAI for software development. (Figures 7 and 8 show respondents’ answers to nine five-point Likert-scale (Oppenheim, 2000) questions.) Of the 554 respondents who indicated a familiarity with AI tools, 205 (37%) somewhat agreed with the view that current litigation will result in limitations placed on using GenAI models for software development, with 76 (14%) respondents strongly agreeing. An additional 109 (20%) respondents were neutral, and the third largest group of 104 developers (19%) somewhat disagreed. The smallest group, 60 developers (11%), strongly disagreed.
While more respondents believed that litigation would result in limitations on the use of GenAI, developers who believed that litigation would not result in limitations were more vocal about their views in their responses to open-ended follow-up questions. Indeed, many respondents indicated that technology would outpace attempts to regulate it, whether through litigation or otherwise. $R_{98}$ commented, “I don’t see any limitations coming due to ongoing litigation—yes, it may cause issues for OpenAI and StabilityAI, but you can’t put the genie back in the lamp.” $R_{507}$ further elaborated: “Pandora’s box has been opened. Even if the current models were legally limited, good luck proving that certain code has been used as training data, and good luck stopping people from running Llama instances locally.” $R_{54}$ acknowledged the existence of legal issues but believed that “the momentum that’s already been built by OpenAI and all the hype around the tools means there will not be the political capital to get any restrictive regulations in place.”
@@
Finally, two respondents expressed that they thought it would not be technology that changed to adapt to the law but rather the law that would change to adapt to technology. $R_{376}$ commented, “The result […] will probably not be limitations on the technology since open-source is rapidly developing and cannot be genuinely regulated in this fashion. Rather, intellectual property, artistic rights, etc. will need to be recalibrated.” $R_{182}$ offered an aspirational approach, stating that they “don’t think it is wise to approach the question of copyright of AI-generated/assisted code as a matter of whether the generated code is or is not copyrightable based primarily on what has been settled. Rather, we should think about what we want it to be, keeping in mind the benefits copyright is intended to bring to humanity.”
@@
5.2. Using one’s code in models’ training data or as output
If regulation of GenAI occurs, whether through legislation or litigation, stakeholders’ perceptions and practices, including software developers, should ideally be considered. Our study, therefore, sought to learn respondents’ views regarding the use of code as training data for generative models.